OpenStudy (anonymous):

You must show your work on each of the following questions. Solve 5x2 = –45x.

OpenStudy (callisto):

\[5x^2 = –45x\]Add 45x to both sides\[5x^2+45x=0\]Factorize the expression. Can you do it?

OpenStudy (poopsiedoodle):

But there's only 1.

OpenStudy (anonymous):

no im lost by the factoring part

OpenStudy (callisto):

Example: \[2a^3 - 2a\]1. factor by grouping\[=2a(a^2-1)\]2. use the identity a^2-b^2 = (a+b)(a-b)\[=2a(a-1)(a+1)\] It's your turn to try. Factorize \(5x^2 +45x\)

OpenStudy (anonymous):

5x2+45x = 5x(x5-1) am i getting it so far?

OpenStudy (anonymous):

?? @callistro

OpenStudy (hba):

What are you doing lol ?

OpenStudy (poopsiedoodle):

\(\Huge\text{Lolwut.}\)

OpenStudy (anonymous):

i need to solve for 5x2 = –45x.

OpenStudy (callisto):

Nope. \[5x^2+45x = 5(x)(x) + 5(9)(x)\]What are the common factors of these two terms?

OpenStudy (anonymous):

not sure

OpenStudy (callisto):

What are common there? :S

OpenStudy (poopsiedoodle):

\[\Huge{x}.\]

OpenStudy (hba):

Lets see it like this 5x^2+45x=0 Now we have to take common you took the common right there which was 5x but you did not leave the rest right. Now try taking common again out of it.

OpenStudy (anonymous):

you mean take it out??

OpenStudy (hba):

Yeah Right @daze I know you can do it.

OpenStudy (hba):

Well if you do not understand this, we can use another method ?

OpenStudy (hba):

@daze Please respond.

OpenStudy (poopsiedoodle):

\(\Huge\text{that'd help.}\)

OpenStudy (anonymous):

well yea could we use another method please

OpenStudy (poopsiedoodle):

\(\huge\text{there we go.}\)

OpenStudy (hba):

Well there are two more methods : One of them is using the quad equation \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Well just compare its general form with your question, Which is, \[ax^2+bx+c=0\] You can now solve it i guess.

OpenStudy (hba):

5x^2+45x=0 Try to simplify your equation first by taking 5 common 5(x^2+9x)=0 x^2+9x+0=0 where a=1,b=9 and c=0 Now put it in quad equation.

OpenStudy (hba):

@daze Now please solve it.